# E-Calib: A Fast, Robust and Accurate Calibration Toolbox for Event Cameras

### v1

@@ IV-B Modified Hierarichal Clustering
The synchronized cylinder centroids $\mathbf{U}_{j}=[u_{j},v_{j}]$ include the actual centers of the pattern circles in addition to false detections of the background clutter, see Fig. 7. This occurs when the sensor observes the scene background as it sweeps the calibration pattern during the data collection step. Hence, a subset of $\mathbf{U}_{j}$, $\mathbf{U}_{i}$ of $M$ elements corresponding to the total number of pattern circles, needs to be extracted to be fed the calibration optimization. We devise a modified hierarchical clustering [40] algorithm, outlined in algorithm 2, that robustly identifies $\mathbf{U}_{i}$ from $\mathbf{U}_{j}$ as the desired image points for calibration optimization and the rest of the $J-M$ centroids as outliers, where $J>M$.
We create a single linkage per each synchronized center to the centroid with the nearest relative euclidean distance $d_{j}$ with its index as
$j\prime^{*}=\operatorname*{arg\,min}_{j\prime}\|(u_{j},v_{j})-d_{1,2,...J\not=j}\|^{2},\$ (8)
@@
This forms the set $\mathbf{D}=\{d_{1},d_{2},\text{ }\cdots,d_{J}\}$ and the corresponding cylinder radii set as $\mathbf{r}=\{r_{1},r_{2},...,r_{J}\}$. Due to the uniform diagonal spacings across the circles and their identical radii, elements of $\mathbf{D}$ and $\mathbf{r}$ corresponding to the pattern circles are very close together compared to the background which are random and sparse, see Fig. 9. In addition, typical camera calibration procedures involve known number of pattern circles, where the maximum cluster cut-off method [40] is utilized to match the detected points to the calibration grid. The cut-off threshold is equal to $M$ and $\mathbf{U}_{i}$ corresponds to the $M$ elements with the minimum variance of $\mathbf{D}$ and $\mathbf{r}$. Therefore, the variance of all unique combinations of sets $\mathbf{D}$ and $\mathbf{r}$ needs to be evaluated such that the indices associated with the minimum variance correspond to $\mathbf{U}_{i}$. We define an index set $\Phi\subset\{1,...,J\}$ such that $|\Phi|=M$ used as follows
$\mathbf{X}_{\Phi}=\{x_{j}\in\mathbf{X}\,|\,j\in\Phi\}.\$ (10)
@@
2:      $\mathbf{U}_{j}=[u_{j},v_{j}]$, $\mathbf{r}=\{r_{1},...,r_{j}\}$
3:Outputs: Circle Centers: $\mathbf{U}_{i}$
4:Find linkages $*j\prime^{*}$, $d_{j}$ using Eq. 8 and 9
@@
This is also understood as extracting all unique combinations of $M$ elements from $\mathbf{X}$, which will be utilized to extract subsets from $\mathbf{D}$ and $\mathbf{r}$. Consequently, $\mathbf{U}_{i}$ is found as
@@
where $\sigma$ is the standard deviation of the set. Note that Eq. 11 is evaluated ${}_{J}C_{M}$ times and $\mathbf{U}_{i}$ is identified with index set $\Phi$ that corresponds to the minimum joint variance norm $\varphi$ of $\mathbf{D}$ and $\mathbf{r}$.
Once $\mathbf{U}_{i}$ is identified for sufficient number of pattern detections $N$, it is utilized as the control points for the calibration optimization, where the estimated parameters are the intrinsic matrix $K$ and the distortion coefficients $\Psi$. Given their initial guess, $\mathbf{U}_{i}$ is shifted by the radial distortion coefficients as
@@
where $N$ is the total number of pattern detections, $M$ is the total number of circles in the calibration pattern, ${}^{P}\mathbf{U}$ are the coordinates of the calibration targets defined in the pattern reference frame, and $\pi$ is the projection of ${}^{P}\mathbf{U}$ into the event camera frame, which is obtained by the camera projection matrix [41]. Eq. 16 is minimized by the levenberg-Marquardt algorithm to obtain $K$ and $\Psi$ [41].
@@ V Experiments
TABLE I: Detection success rate and reprojection error on the ECam_ACircles dataset. Notice that high detection success rates and low reprojection errors are maintained even at challenging lighting conditions demonstrating robustness against noise.
Metric Dataset
davis_3x7
gdlight
davis_3x7
lowlight
davis_3x9
gdlight
davis_3x9
lowlight
davis_4x11
gdlight
davis_4x11
lowlight
dvx_3x7
gdlight
dvx_3x9
gdlight
dvx_4x11
gdlight
Data Sequence
Time (s)
17.04 15.37 27.02 24.04 19.07 18.53 15.53 27.04 14.99
Total
Detections
438 362 709 648 437 392 330 672 312
Detection
Success Rate (%)
84.12 80.44 90.89 89.99 76.84 71.68 72.34 82.98 74.44
$\mathbf{\zeta_{r}}$ (Px)
0.13 0.19 0.13 0.17 0.16 0.21 0.46 0.51 0.43
@@ V-A Experimental Setup
We have tested the proposed algorithm in various challenging scenarios demonstrated in our ECam_ACircles dataset. To show that our proposed approach is resolution agnostic, event camera models with different resolutions were calibrated by the devised algorithm; the DAVIS346 $(346\times 260)$ and DVXplorer ($640\times 480$), see Fig. 10. Both cameras are interfaced with robot operating system (ROS) using a USB 3.0 terminal on the host computer. We have also tested our method on three calibration pattern variants of ECam_ACircles, of dimensions $3\times 7$ and 34 mm diagonal spacing, $3\times 9$ and 27 mm diagonal spacing, and $4\times 11$ and 24 mm diagonal spacing. This is of paramount importance to prove that the developed calibration framework successfully calibrates event cameras regardless of the pattern geometric properties. Finally, the work of the paper was assessed in good (93.21 Lux) and poorly lighted environments (8.72 Lux) to demonstrate robustness against illumination variation and increasing noise with degraded lighting conditions. Videos of the experiments are available through the following link: https://youtu.be/4giQn6rt-48.
We evaluated our method on the aforementioned experiments based on three metrics: 1) Detection success rate, 2) calibration reprojection error, and 3) positioning error of the estimated extrinsics. While comparing the calibration reprojection error and the obtained intrinsics to the DAVIS346 frames can be sufficient for validation, inaccurate intrinsics can be obtained as the solver can converge to a local minimum, even though a low reprojection error is achieved. This is usually witnessed with the lens distortion coefficients. Thus, we additionally compare the calibrated extrinsics to the ground truth pose of the camera, which is provided by OptiTrack Prime 13 motion capture system to further evaluate our method fairly. It is worth mentioning that the positioning measurements from OptiTrack and the estimated extrinsics are synchronized using soft time synchronization [42].
@@ V-B Calibration Accuracy Evaluation
To obtain optimal intrinsics of the event camera, 1) sufficient number of detections of the asymmetric circle grid is required and 2) the detected circle centers need to be localized with sub-pixel accuracy in the image plane. Accordingly, we compare our eRWLS feature extraction method to the detection performance of MATLAB’s state-of-the-art detector, detectCircleGridPoints [43], applied on the frames of the DAVIS APS sensor and MR algorithm for $davis\_4x11\_gdlight$ and $dvx\_4x11\_dataset$ datasets, demonstrated in Fig. 11. In addition, the detection success rate evaluates the performance of our modified hierarchical clustering algorithm to detect the circle grid even in the presence of scene background. detectCircleGridPoints fails to detect the calibration pattern on the DAVIS frames due to the fact that motion blur is witnessed and the features of the calibration grid are degraded. On the other hand, even though the MR frames suffer from noise and artifacts significantly degrading the image quality and the calibration grid is not extracted.
Figure 12: Reprojection errors on $davis\_3x9\_gdlight$ dataset where $\zeta_{r}=0.13$ pixels was obtained.
@@
To the contrary, our calibration tool, leveraging the asynchronous nature of events, detects the calibration pattern even when high speed motion is induced. We first evaluate the proposed calibration approach in terms of the detection success rate on ECam_ACircles dataset to ensure that calibration is attainable, where the detection success rate is defined as the fraction of times the pattern was detected against the total number of detections that can be obtained. It is worth noting that the total number of detections of the calibration pattern is equal to the data sequence time divided by the time step between the spatiotemporal windows, which we have defined as 33 ms. The detection success rates are shown in Table I. Notice that the detection success rate is high for both sensors showing that our approach proves to be resolution agnostic. In addition, the detection success rate is maintained even at challenging illumination conditions demonstrating that the proposed eRWLS is robust against noise and the modified hierarchical clustering algorithm identifies the calibration pattern despite the presence of the noisy background.
@@
Figure 14: Data sequences on the $davis\_4x11\_gdlight$ (top) and $dvx\_4x11\_dataset$ (bottom), where state-of-the-art methods fail to detect the calibration pattern, while our method robustly detects the asymmetric circle grid.
Having a sufficient number of pattern detections, the calibration accuracy is mainly dependent on the feature extraction performance of eRWLS since a state-of-the-art calibration optimization is utilized to obtain the sensor parameters. Hence, the calibration performance is evaluated in terms of the reprojection error and the deviation of the obtained intrinsic parameters from the conventional frames. The reprojection errors $\zeta_{r}$, evaluated similar to [41], are reported in Table I and Fig. 12, while the obtained intrinsics compared against the conventional frames is reported in Table II. Note that this comparison is only applicable on $davis\_3x9\_gdlight$ because calibration was only successful using the DAVIS frames on the aforementioned dataset due to the slow sensor motion and adequate lighting conditions. The obtained results show that eRWLS extracts the pattern circle centers with the required accuracy. Moreover, low reprojection errors are maintained even in poorly illuminated environments, where sensor noise drastically increases. This shows the importance of time synchronization of the predicted circle centers for maintaining optimal calibration performance.
@@
TABLE II: The obtained intrinsics of the event camera by our method are compared against the parameters acquired by the DAVIS frames on $davis\_3x9\_gdlight$.
@@
TABLE III: Detection success rate and reprojection error on the ECam_ACircles dataset. Notice that high detection success rates and low reprojection errors are maintained even at challenging lighting conditions demonstrating robustness against noise.
@@
TABLE IV: The proposed event-based calibration algorithm is benchmarked against the works of Reinbacher et al [44], Muglikar et al. [17], Scheerlinck et al. [32] and Huang et al. [19] on the $davis\_3x9\_gdlight$. The dash indicates the metric is not applicable while † represents failure of obtaining the parameter.

### v2

@@ IV-B Modified Hierarichal Clustering
The synchronized cylinder centers $\mathbf{U}=\{(u_{j},v_{j})\}_{j=1}^{J}$ include the actual centers of the pattern circles in addition to false detections of the background clutter, see Fig. 7. This occurs when the sensor observes the scene background as it sweeps the calibration pattern during the data collection step. Hence, a subset of $\mathbf{U}$, $\mathbf{U}_{M}\subset\mathbf{U}$ of $M$ elements corresponding to the total number of pattern circles, needs to be extracted to be fed the calibration optimization. We devise a modified hierarchical clustering [37] algorithm, outlined in algorithm 2, that robustly identifies $\mathbf{U}_{M}$ from $\mathbf{U}$ as the desired image points for calibration optimization and the rest of the $J-M$ centers as outliers, where $J>M$. It is worth mentioning that the clustering algorithm in [37] is utilized and we rather modify the input such that it is represented by the uniform geometrical representation of the circles in the calibration pattern to distinguish it from the background clutter.
We create a single linkage per each synchronized center to the centers with the nearest relative euclidean distance $d_{j}$ with its index as
$j\prime^{*}=\operatorname*{arg\,min}_{j\prime=1,2,...J\not=j}\|(u_{j},v_{j})-(u_{j\prime},v_{j\prime})\|^{2},\$ (8)
@@
This forms the set $\mathbf{D}=\{d_{1},d_{2},\text{ }\cdots,d_{J}\}$ and the corresponding cylinder radii set as $\mathbf{r}=\{r_{1},r_{2},...,r_{J}\}$. Due to the uniform diagonal spacings across the circles and their identical radii, elements of $\mathbf{D}$ and $\mathbf{r}$ corresponding to the pattern circles are very close together compared to the background which are random and sparse, see Fig. 9. In addition, typical camera calibration procedures involve known number of pattern circles, where the maximum cluster cut-off method [37] is utilized to match the detected points to the calibration grid. The cut-off threshold is equal to $M$ and $\mathbf{U}_{M}$ corresponds to the $M$ elements with the minimum variance of $\mathbf{D}$ and $\mathbf{r}$. Therefore, the variance of all unique combinations of sets $\mathbf{D}$ and $\mathbf{r}$ needs to be evaluated such that the indices associated with the minimum variance correspond to $\mathbf{U}_{M}$. We define an index set $\Phi\subset\{1,...,J\}$ such that $|\Phi|=M$ used as follows
$\mathbf{S}_{\Phi}=\{s_{j}\in\mathbf{S}\,|\,j\in\Phi\}.\$ (10)
@@
2:      $\mathbf{U}_{j}=\{(u_{j},v_{j})\}_{j=1}^{J}$, $\mathbf{r}=\{r_{1},...,r_{j}\}$
3:Outputs: Pattern Circle Centers: $\mathbf{U}_{M}$
4:Find linkages $j\prime^{*}$, $d_{j}$ using Eq. 8 and 9
@@
This is also understood as extracting all unique combinations of $M$ elements from $\mathbf{S}$, which will be utilized to extract subsets from $\mathbf{D}$ and $\mathbf{r}$. Consequently, $\mathbf{U}_{M}$ is found as
@@
where $\sigma$ is the standard deviation of the set. Note that Eq. 11 is evaluated ${}_{J}C_{M}$ times and $\mathbf{U}_{M}$ is identified with index set $\Phi$ that corresponds to the minimum joint variance norm $\varphi$ of $\mathbf{D}$ and $\mathbf{r}$.
Once $\mathbf{U}_{M}$ is identified for sufficient number of pattern detections $N$, it is utilized as the control points for the calibration optimization, where the estimated parameters are the intrinsic matrix $K$ and the distortion coefficients $\Psi$. Given their initial guess, $\mathbf{U}_{M}$ is shifted by the radial distortion coefficients as
@@
where $N$ is the total number of pattern detections, $M$ is the total number of circles in the calibration pattern, ${}^{P}\mathbf{U}$ are the coordinates of the calibration targets defined in the pattern reference frame, and $\pi$ is the projection of ${}^{P}\mathbf{U}$ into the event camera frame, which is obtained by the camera projection matrix defined in Eq. 2. Eq. 16 is minimized by the levenberg-Marquardt algorithm to obtain $K$ and $\Psi$ [41].
@@ V-A Experimental Setup
We have tested the proposed algorithm in various challenging scenarios demonstrated in our ECam_ACircles dataset. To show that our proposed approach is resolution agnostic, event camera models with different resolutions were calibrated by the devised algorithm; the DAVIS346 $(346\times 260)$ and DVXplorer ($640\times 480$), see Fig. 10. Both cameras are interfaced with robot operating system (ROS) using a USB 3.0 terminal on the host computer. We have also tested our method on three calibration pattern variants of ECam_ACircles, of dimensions $3\times 7$ and 34 mm diagonal spacing, $3\times 9$ and 27 mm diagonal spacing, and $4\times 11$ and 24 mm diagonal spacing. This is of paramount importance to prove that the developed calibration framework successfully calibrates event cameras regardless of the pattern geometric properties. In addition, the work of the paper was assessed in good (93.21 Lux) and poorly lighted environments (8.72 Lux) to demonstrate robustness against illumination variation and increasing noise with degraded lighting conditions. Finally, E-Calib was validated on three different paths, i.e. cone, square, and spiral, at varying speeds to test the method against harsh motion trajectories including out-of-focus scenes. Videos of the experiments are available through the following link: https://youtu.be/4giQn6rt-48.
We evaluated our method on the aforementioned experiments based on three metrics: 1) Detection success rate, 2) calibration reprojection error, and 3) positioning error of the estimated extrinsics. While comparing the calibration reprojection error and the obtained intrinsics to the DAVIS346 frames can be sufficient for validation, inaccurate intrinsics can be obtained as the solver can converge to a local minimum, even though a low reprojection error is achieved. This is usually witnessed with the lens distortion coefficients. Thus, we additionally compare the calibrated extrinsics to the ground truth pose of the camera, which is provided by OptiTrack Prime 13 motion capture system to further evaluate our method fairly. It is worth mentioning that the positioning measurements from OptiTrack and the estimated extrinsics are synchronized using soft time synchronization [42]. In addition, the calibration patterns coordinates were registered to the OptiTrack frame by extrinsic calibration using the method in [43, 7].
Figure 12: Reprojection errors on davis_3x9_cone_gdlight dataset where $\zeta_{r}=0.13$ pixels was obtained.
TABLE I: Detection success rate and reprojection error on the cone trajectory sequences of the ECam_ACircles dataset.
Notice that high detection success rates and low reprojection errors are maintained even at challenging lighting conditions demonstrating robustness against noise. Metric Dataset davis_3x7 gdlight davis_3x7 lowlight davis_3x9 gdlight davis_3x9 lowlight davis_4x11 gdlight davis_4x11 lowlight dvx_3x7 gdlight dvx_3x9 gdlight dvx_4x11 gdlight Data Sequence Time (s) 17.04 15.37 27.02 24.04 19.07 18.53 15.53 27.04 14.99 Total Detections 438 362 709 648 437 392 330 672 312 Detection Success Rate (%) 84.12 80.44 90.89 89.99 76.84 71.68 72.34 82.98 74.44 $\mathbf{\zeta_{r}}$ (Px) 0.13 0.19 0.13 0.17 0.16 0.21 0.46 0.51 0.43
TABLE II: Calibration accuracy of of E-Calib eRWLS compared against OLS, MLESAC [45], and RANSAC [46], in terms of the reprojection error $\zeta_{r}$.
Method Dataset
davis_3x7_cone
gdlight
davis_3x7_cone
lowlight
davis_3x9_cone
gdlight
davis_3x9_cone
lowlight
davis_4x11_cone
gdlight
davis_4x11_cone
lowlight
OLS 0.39 0.69 0.40 0.53 0.43 0.49
MLESAC 0.43 0.48 0.37 0.36 0.33 0.37
Hough Transform
with RANSAC
0.36 0.37 0.29 0.31 0.36 0.41
eRWLS 0.13 0.19 0.13 0.17 0.16 0.21
@@ V-B Calibration Accuracy Evaluation
To obtain optimal intrinsics of the event camera, 1) sufficient number of detections of the asymmetric circle grid is required and 2) the detected circle centers need to be localized with sub-pixel accuracy in the image plane. Accordingly, we compare our eRWLS feature extraction method to the detection performance of MATLAB’s state-of-the-art detector, detectCircleGridPoints [44], applied on the frames of the DAVIS APS sensor and MR algorithm for davis_4x11_cone_gdlight and dvx_4x11_cone_gdlight datasets, demonstrated in Fig. 11. In addition, the detection success rate evaluates the performance of our modified hierarchical clustering algorithm to detect the circle grid even in the presence of scene background. detectCircleGridPoints fails to detect the calibration pattern on the DAVIS frames due to the fact that motion blur is witnessed and the features of the calibration grid are degraded. On the other hand, even though the MR frames suffer from noise and artifacts significantly degrading the image quality and the calibration grid is not extracted.
@@
To the contrary, our calibration tool, leveraging the asynchronous nature of events, detects the calibration pattern even when high speed motion is induced. We first evaluate the proposed calibration approach in terms of the detection success rate on ECam_ACircles dataset to ensure that calibration is attainable, where the detection success rate is defined as the fraction of times the pattern was detected against the total number of detections that can be obtained. It is worth noting that the total number of detections of the calibration pattern is equal to the data sequence time divided by the time step between the spatiotemporal windows, which we have defined as 33 ms. The detection success rates are shown in Table I on the cone trajectory data sequences of the ECam_ACircles dataset. We also report the mean reprojection error for all DAVIS and DVXplorer data sequences in Table VI, where a mean reprojection error of 0.16 Px and 0.52 Px is obtained for DAVIS and DVXplorer calibrations, respectively. Notice that the detection success rate is high for both sensors showing that our approach proves to be resolution agnostic. In addition, the detection success rate is maintained even at challenging illumination conditions demonstrating that the proposed eRWLS is robust against noise and the modified hierarchical clustering algorithm identifies the calibration pattern despite the presence of the noisy background.
@@
Figure 14: Data sequences on the davis_4x11_cone_gdlight (top) and dvx_4x11_cone_gdlight (bottom), where state-of-the-art methods fail to detect the calibration pattern, while our method robustly detects the asymmetric circle grid.
@@
TABLE III: The obtained intrinsics of the event camera by our method are compared against the parameters acquired by the DAVIS frames on davis_3x9_cone_gdlight.
@@
Having a sufficient number of pattern detections, the calibration accuracy is mainly dependent on the feature extraction performance of eRWLS since a state-of-the-art calibration optimization is utilized to obtain the sensor parameters. Hence, the calibration performance is evaluated in terms of the reprojection error and the deviation of the obtained intrinsic parameters from the conventional frames. The reprojection errors $\zeta_{r}$, evaluated similar to [41], are reported in Table I and Fig. 12, while the obtained intrinsics compared against the conventional frames is reported in Table III. Note that this comparison is only applicable on davis_3x9_cone_gdlight because calibration was only successful using the DAVIS frames on the aforementioned dataset due to the slow sensor motion and adequate lighting conditions. The obtained results show that eRWLS extracts the pattern circle centers with the required accuracy. Moreover, low reprojection errors are maintained even in poorly illuminated environments, where sensor noise drastically increases.
We also motivate the use of eRWLS by comparing its performance to cylinder fitting using OLS, in terms of the reprojection error. In addition, we also provide comparisons with respect to state-of-the-art cylinder detection approaches, namely maximum-likelihood estimation by random sampling consensus (MLESAC) [45] and random sample consensus (RANSAC) [47, 46] with Hough Transform, reported in Table II on the DAVIS cone path data sequence. Notice that OLS performs poorly in poor-lit conditions. On the other hand, MLESAC and RANSAC show consistent performance in poor in poorly lighted environments, still E-Calib outperforms these approaches in terms of the reprojection error $\zeta_{r}$, ensuring high calibration accuracy for event cameras.
TABLE IV: Pose errors on cone trajectory data sequences of the ECam_ACircles dataset. Notice that high detection success rates and low reprojection errors are maintained even at challenging lighting conditions demonstrating robustness against noise, which is quite high in low illumination conditions.
@@
Finally, as mentioned in section III-B, we have formulated spatiotemporal windows based on 4000 events. Even though though the number of events is set as a hyperparameter, we study the effect of the number of events on the calibration accuracy. Large number of events can lead to rather curved cylinders instead of slanted cylinders due to varying optical flow. Accordingly, we evaluate different number of events when formulating spatiotemporal windows and validate eRWLS feature extraction performance in terms of the reprojection error and detection success rate when curved cylinders are present, see Table V reporting the calibration accuracy on dvx_3x9_cone_gdlight. Notice that the reprojection error is relatively consistent, even when large number of events, i.e. 15,000, leading to curved cylinders to be present. This demonstrates the robustness of eRWLS with the weighing factor that penalizes outliers, considered as points along the curved cylinder centerline.
TABLE V: Effect of the number of events on the eRWLS feature extraction performance on the dvx_3x9_cone_gdlight data sequence.
No. Events $\zeta_{r}$ (Px)
Detection
Success Rate (%)
4000 0.46 72.34
10000 0.49 70.21
15000 0.54 74.33
25000 0.47 77.21
TABLE VI: Mean reprojection errors $\zeta_{r}$ obtained by E-Calib on all ECam_ACircles calibration trajectories.
Data Sequence
Trajectory
DAVIS DVXplorer
Cone 0.14 0.47
Square 0.11 0.46
Spiral 0.18 0.54
Out-of-Focus
Scenes
0.21 0.59
TABLE VII: Mean translation $\hat{\eta}_{t}$ (cm) and rotation errors $\hat{\eta}_{r}$ $({}^{\circ})$ obtained by E-Calib on all ECam_ACircles calibration trajectories.
Data Sequence Trajectory DAVIS DVXplorer
$\hat{\eta}_{t}$ (cm) $\hat{\eta}_{r}$ () $\hat{\eta}_{t}$ (cm) $\hat{\eta}_{r}$ ()
Cone 0.916 0.930 0.642 0.735
Square 0.879 1.021 0.777 0.822
Spiral 0.992 0.991 0.719 0.801
Out-of-Focus
Scenes
1.612 1.176 1.113 0.939
TABLE VIII: The proposed event-based calibration algorithm is benchmarked against the works of Reinbacher et al [48], Muglikar et al. [17], Scheerlinck et al. [32] and Huang et al. [19] on the davis_3x9_cone_gdlight. The dash indicates the metric is not applicable while † represents failure of obtaining the parameter.